# Species Quantum Mechanics

### v2

@@ Abstract
In this note we introduce some concepts of Species Quantum Mechanics. Specifically, we consider quantum operators that correspond to the species number $N_{s}$ and the tower mass scale $m_{t}$ in the context of the swampland distance conjecture. We discuss the commutation relations, a possible wave function, and symplectic duality transformations on the conjugate variables. Furthermore, we argue that the Castellano-Ruiz-Valenzuela (CRV) pattern is a consequence of the canonical commutation rules of moduli space quantum mechanics. We also connect the canonical quantization to the periods of ${\cal N}=2$ Calabi-Yau compactifications to explore other aspects of the CRV pattern, including its possible connection to the Ooguri-Vafa-Verlinde black hole quantization procedure.
@@ II.1 Canonical variables and commutation relations
We will now treat the scalar fields $t^{i}$ as quantum fields, and we first want to determine their conjugate momenta $\pi_{t}$. Following the standard canonical quantization procedure, the conjugate momenta in quantum field theory are given by $\pi_{i,\mu}={\delta{\cal L}_{\rm EFT}\over\delta(\partial_{\mu}t_{i})}=g_{ij}\ \partial_{\mu}t^{j}$. But to obtain a sensible quantization, i.e. in order to describe a quantum mechanical system, the canonical momenta should only be defined w.r.t to the time derivative. Therefore, as also considered in the context of finiteness Delgado:2024skw , we assume that all spatial coordinates are compactified and only the time coordinate $\tau$ is left. Then following the standard canonical quantization procedure, the canonical momenta are just given by the time-derivative of the moduli fields:
@@
The commutator (22) leads to a Heisenberg uncertainty relation of the form
@@
The interpretation of this Heisenberg uncertainty is very interesting: it means that if one does a precise measurement of the number of species, or respectively the species scale, quantum mechanics forbids to measure the time variation of the tower mass scale with high precision. Of course the reverse statement also is true.
@@ II.2 The CRV pattern
It corresponds to a Heisenberg uncertainty relation of the following form
@@ II.4 Species wave function
The wave function $\psi_{s}(t)$ of species quantum mechanics should correspond to some “wave function of the universe,” being the solution of the Schrödinger equation of the form
@@
with the Hamiltonian as given in (8). [^3] Some ideas about the wave function of the universe in brane world models are given in Anchordoqui:2000du . The solutions of the Schrödinger equation will provide for example the expectation value of the number of species: $\langle N_{s}\rangle=\langle\psi_{s}|N_{s}|\psi_{s}\rangle$.
@@ III.1 Type IIA Calabi-Yau compactifications and cubic prepotential
Asking for the volumes of all complex submanifolds of the CY to be positive defines a region of the moduli space known as Kähler cone. For the supergravity approximation to be reliable, one should remain within the bulk of this region.
@@ IV Summary
This leads to an uncertainty relation basically saying that a large species uncertainty implies a small uncertainty in the time variation of the tower mass and vice versa. As a consequence, we also demonstrated that the CRV pattern is closely related to this pair of conjugate variables. Furthermore we also investigated the corresponding canonical commutation relations of Calabi-Yau period vectors in ${\cal N}=2$ string compactifications. Here we highlighted a universal relation of the species number $N_{s}$ and the tower mass scale $m_{t}$ in terms of the genus zero and the genus one prepotentials ${\cal F}_{0}$ and ${\cal F}_{1}$, which led us to a proposal for the CRV pattern including instanton corrections.
@@
This wave function corresponds to a kind of mini-superspace wave function of quantum gravity. It would be interested to see, which expectation values for $N_{s}$ and $m_{t}$ have high or low probability (see also LopesCardoso:2006efn ; LopesCardoso:2006xue ). Furthermore, we noted that the species wave function $\psi_{s}$ is very similar to the OVV wave function in ${\cal N}=2$ string compactifications. We also discussed some cases with non-vanishing potentials like those characterizing the dark dimension scenario.

### v3

@@ Abstract
In this note we take the initial steps towards a formulation of Species Quantum Mechanics. Specifically, we consider quantum operators that correspond to the species number $N_{s}$ and the tower mass scale $m_{t}$ in the context of the swampland distance conjecture. We discuss the commutation relations, a possible wave function, and symplectic duality transformations on the conjugate variables. Furthermore, we argue that the Castellano-Ruiz-Valenzuela (CRV) pattern is a consequence of the canonical commutation rules of moduli space quantum mechanics. We also connect the canonical quantization to the periods of ${\cal N}=2$ Calabi-Yau compactifications to explore other aspects of the CRV pattern, including its possible connection to the Ooguri-Vafa-Verlinde black hole quantization procedure.
@@ II.1 Canonical variables and commutation relations
We will now treat the scalar fields $t^{i}$ as quantum fields, and we first want to determine their conjugate momenta $\pi_{t}$. Following the standard canonical quantization procedure, the conjugate momenta in quantum field theory are given by $\pi_{i,\mu}={\delta{\cal L}_{\rm EFT}\over\delta(\partial_{\mu}t_{i})}=g_{ij}\ \partial_{\mu}t^{j}$. But to obtain a sensible quantization, i.e. in order to describe a quantum mechanical system, the canonical momenta should only be defined w.r.t to the time derivative. Therefore, as also considered in the context of finiteness Delgado:2024skw , we assume that all spatial coordinates are compactified and only the time coordinate $\tau$ is left. [^2] Actually, in a way we do not really compactified to one dimension, but we just ignore the three space-like dimensions. A subtle point of this truncation is that we work in four-dimensional Planck units and so the relations between the tower masses and the species scale derived in the four-dimensional EFT do not change. Then following the standard canonical quantization procedure, the canonical momenta are just given by the time-derivative of the moduli fields:
@@
The commutator (22) leads to a Robertson uncertainty relation of the form
@@
The interpretation of this Robertson uncertainty is very interesting: it means that if one does a precise measurement of the number of species, or respectively the species scale, quantum mechanics forbids to measure the time variation of the tower mass scale with high precision. Of course the reverse statement also is true.
@@ II.2 The CRV pattern
It corresponds to a Robertson uncertainty relation of the following form
@@ II.4 Species wave function
It has long been known that the relevant field equation in quantum gravity is the Wheeler-DeWitt equation (WDW) Wheeler:1968iap ; DeWitt:1967yk . More concretely, in the Arnowitt–Deser–Misner (a.k.a. ADM) formalism Arnowitt:1959ah the WDW equation takes the form
${\cal H}\Psi[h]=0\,$, (37)
where $h$ is the space-time metric and the Hamiltonian is given by
${\cal H}={1\over\sqrt{h}}\left(\pi^{ij}\pi_{ij}-{1\over 2}\pi^{2}\right)-\sqrt{h}\ R\,.$ (38)
Note that time and time derivatives do not appear in the WDW equation, i.e. no time evolution emerges in terms of the WDW wave function dependence. [^4] Some ideas about the wave function of the universe in brane world models are given in Anchordoqui:2000du .
[…]
@@
with the Hamiltonian as given in (8). The solutions of the Schrödinger equation will provide the expectations values for observables $\hat{\cal O}(t)$:
$\langle\hat{\cal O}(t)\rangle=\langle\psi_{s}|\hat{\cal O}(t)|\psi_{s}\rangle\,.$ (40)
For example the expectation values of the number of species, $\langle N_{s}\rangle=\langle\psi_{s}|N_{s}|\psi_{s}\rangle$, or of the species scale, $\langle\Lambda_{s}\rangle=\langle\psi_{s}|\Lambda_{s}|\psi_{s}\rangle$, are important observables that can be measured in principle by experiments.
@@
Before proceeding, a brief word of caution is necessary. While the plane wave given in (43) is a definite-momentum solution to the Schrödinger equation, it is not a localized solution. Actually, $\psi_{s}$ as given in (43) is not normalizable over infinite moduli space. Thus, to determine e.g. $\langle N_{s}\rangle$ the wave function $\psi_{s}$ must be a wave packet with a finite extent in moduli space (e.g. a Gaussian wave packet), which is a superposition of different momentum (or wave number $k_{\phi}$) eigenstates,
$\psi_{s}(t)$ $\sim$ $\frac{1}{\sqrt{2\pi}}\int_{-\infty}^{+\infty}\varphi(k_{\phi})\ e^{ik_{\phi}\phi(t)}\ dk_{\phi}$ (47)
$\sim$ $e^{-\frac{[\phi(t)-\phi_{0}]^{2}}{2\sigma^{2}}}\ e^{ik_{\phi_{0}}\phi(t)}\,$,
where $\sigma$ determines the width of the packet and the expectation value $\langle\phi\rangle\sim\phi_{0}$ aligns with the classical trajectory. The expectation value then evolves over time according to Ehrenfest’s theorem, where its velocity is related to the average momentum $\langle\pi_{\phi}\rangle=k_{\phi_{0}}$. Using (15) and (18) it is now straightforward to obtain $\langle N_{s}\rangle$.
@@ III.1 Type IIA Calabi-Yau compactifications and cubic prepotential
Asking for the volumes of all complex submanifolds of the CY to be positive defines a region of the moduli space known as Kähler cone. For the supergravity approximation to be reliable, one should remain within the bulk of this region. It is convenient to pass to the real basis of fields $t^{i}$, since their axionic partners $b^{i}$ do not enter the Kähler potential. Hence, throughout we set $b^{i}=0$.
@@ IV Summary and Conclusions
This leads to an uncertainty relation basically saying that a large species uncertainty implies a small uncertainty in the time variation of the tower mass and vice versa. As a consequence, we also demonstrated that the CRV pattern is closely related to this pair of conjugate variables. [^7] As a side note, it is worth mentioning that the commutator in (28) always equates to ($i$ times) the CRV pattern, because this relation is a proven property of quantum mechanical operators, as demonstrated in (26). Now, assuming that the CRV pattern holds, the operators within the commutator may be treated as conjugates, but of course we can explicitely use the pattern in the asymptotic regime. We have then examined the pattern for three possible types of light towers in infinite-distance limits, where the pattern has been shown to be satisfied. Furthermore we also investigated the corresponding canonical commutation relations of Calabi-Yau period vectors in ${\cal N}=2$ string compactifications. Here we highlighted a universal relation of the species number $N_{s}$ and the tower mass scale $m_{t}$ in terms of the genus zero and the genus one prepotentials ${\cal F}_{0}$ and ${\cal F}_{1}$, which led us to a proposal for the CRV pattern including instanton corrections.
@@
This wave function corresponds to a kind of mini-superspace wave function of quantum gravity. It is important to stress that herein we do not consider the quantum gravity over the four-dimensional space-time, but of the scalars that come from the compact part of the metric. It would be interested to see, which expectation values for $N_{s}$ and $m_{t}$ have high or low probability (see also LopesCardoso:2006efn ; LopesCardoso:2006xue ). Furthermore, we noted that the species wave function $\psi_{s}$ is very similar to the OVV wave function in ${\cal N}=2$ string compactifications. We also discussed some cases with non-vanishing potentials like those characterizing the dark dimension scenario.